# Bounds and Constructions of High-Memory Spatially-Coupled Codes

### v1

@@ I Introduction
Spatially-coupled low-density parity-check (SC-LDPC) codes have attracted much attention due to their threshold saturation property [1, 2, 3] and the capacity-achieving performance over general binary memoryless channel [4]. In addition, the highly regular structure makes it possible to decode using low-latency windowed decoding (WD) [5], [6]. Research on SC-LDPC codes can be traced back to the convolutional LDPC codes introduced by Felström and Zigangirov [7].
In the practical data storage systems, the frame error rate (FER) requirement is below $10^{-12}$ [8, 9, 10]. Therefore, it is crucial to construct SC-LDPC codes with good bit-error rate (BER) performance in both waterfall and error floor regions under belief-propagation (BP) decoders. In the low-SNR region, the performance of the SC-LDPC codes are mainly determined by their asymptotic properties (such as decoding threshold) [4], and density evolution (DE) techniques can be used to analyze the decoding threshold of the code. In the high-SNR region, the performance of SC-LDPC codes is determined by finite-length characteristics associated with harmful topological structures (e.g. cycles and trapping sets) [9], [10].
@@
In this paper, we consider the non-uniform protograph-based SC-LDPC codes in the same probability metric as [14], and provide mathematical support for the construction of SC-LDPC codes through probabilistic combinatorial tools.
@@ II-A QC-SC-LDPC Codes
In this section, we introduce the construction process of QC-SC-LDPC codes using QC-LDPC as underlying block codes. The QC structure not only enables efficient encoding and decoding implementation to SC-LDPC codes [15], [16], but also makes the constructed codes exhibit anytime-reliable properties [17], [18].
We consider the QC-SC-LDPC code of Type-I. First, we start with an all-one $(\gamma,\kappa)$ base matrix $\mathbf{H}$, which corresponds to a fully connected bipartite base graph. As the name suggests, SC-LDPC codes are generally constructed by coupling a series of identical block codes into a chain. This operation has different names in different perspectives, but the most widely intuitive is the matrix perspective (edge-spreading), which is seen as dividing the edges of the base matrix $\mathbf{H}$ into $m+1$ component matrices $\{\mathbf{H}_{0},...,\mathbf{H}_{m}\}$ of the same size, such that $\mathbf{H}=\sum_{l=0}^{m}\mathbf{H}_{l}$, where $m$ is memory. We record the exact way in which the edge spreads in the $(\gamma,\kappa)$ partition matrix $\mathbf{P}$, which means that $\mathbf{P}(i,j)=k,0\leq k\leq m,i\in[\gamma],j\in[\kappa]$, if and only if $\mathbf{H}_{k}(i,j)=1,\mathbf{H}_{l}(i,j)=0,\forall l\neq k,0\leq l\leq m$. Generally, the base matrices and edge-spreading patterns may differ among uncoupled codes, like those constructed via the random spreading method for the proof of universality . However, considering the implementation complexity in practical systems, the base matrices and the edge-spreading patterns of all uncoupled codes are assumed to be identical in this paper. By concatenating these component matrices vertically from $\mathbf{H}_{0}$ to $\mathbf{H}_{m}$, a replica is obtained. By concatenating $L$ of these identical replicas horizontally in space, we get the protograph matrix $\mathbf{H}_{SC}^{P}$ of the SC-LDPC code, where $L$ is called the coupling length.
To construct QC-SC-LDPC codes, after the edge-spreading process, we perform the lifting operation on the protograph. The lifting for every uncoupled code is identical, which facilitates the practical implementation of the windowed decoder. Let $\mathbf{L}$ be a $\gamma\times\kappa$ matrix called the lifting matrix. $\sigma$ denotes the $Z\times Z$ circulant matrix obtained by cyclically shifting the columns of an identity matrix one unit to the left. For all $i\in[\gamma],j\in[\kappa]$, if $\mathbf{L}(i,j)=x$ where $x\in\{0,\cdots,Z-1\}$, then replace the ones in $\mathbf{H}_{SC}^{P}$ corresponding to $\mathbf{H}(i,j)$ with the $\sigma^{x}$ matrix, and replace the remaining zeros with $Z\times Z$ zero matrices. Under this notation, the parameters of an QC-SC-LDPC code $\mathbf{H}_{SC}$ are described by $(\gamma,\kappa,m,Z,L)$.
@@ II-B Major Harmful Structures
Despite the fact that SC-LDPC codes offer improved performance compared to their underlying block codes, they still experience performance degradation at high signal-to-noise ratios (SNRs) over the additive white Gaussian noise (AWGN) channel under iterative decoding, resulting in an error floor phenomenon. This error floor is primarily caused by small structures in the code’s Tanner graph, known as trapping sets (TSs), which, in turn, consist of small cycles. We review the key definitions:
@@
Let $\mathcal{C}_{2g}$ be the set of cycle-2g candidates in the base matrix, where $g\in\mathbb{N},g\geq 2$. Denote $c_{2g}\in\mathcal{C}_{2g}$ by $(j_{1},i_{1},j_{2},i_{2},\dots,j_{g},i_{g})$, where $(i_{k},j_{k})$ and $(i_{k},j_{k+1})$ for $1\leq k\leq g$, with $j_{g+1}=j_{1}$, are nodes of $c_{2g}$ in $\mathbf{H}$, $\mathbf{P}$, and $\mathbf{L}$. Then $c_{2g}$ becomes a cycle candidate in the protograph after the edge-spreading operation if and only if the following condition holds [19]:
@@
To reduce the complexity of optimizing a QC-SC-LDPC codes, the process is typically divided into two stages. First, matrix $\mathbf{P}$ is optimized to minimize cycle candidates in the protograph. This sets the stage for optimizing matrix $\mathbf{L}$, aiming to further reduce cycles in the Tanner graph. This two-step method efficiently streamlines optimization and reduces cycles, as shown in [12], [14].
In the remainder of this paper, we focus on QC-SC codes for the AWGN channel, where the dominant harmful structures are the low-weight absorbing sets (ASs) [12]. The ASs are defined as follows:
@@ II-C Lovász Local Lemma
The Lovász Local Lemma (LLL), introduced by Erdős and Lovász in 1975, has become one of the most important probabilistic methods. LLL provides sufficient conditions under which a set of undesirable events $\mathcal{A}$ in a probability space $\Omega$ can be simultaneously avoided, meaning the conditions for $P(\cap_{A\in\mathcal{A}}\bar{A})>0$ to hold. The relationships between these bad events are depicted by an undirected graph called the dependency graph $G_{D}=([n],E)$, where each bad event corresponds to a vertex in the dependency graph. An event $A_{i}$ is independent of $\{A_{j}:j\neq i,j\notin\mathcal{N}(i)\}$, where $\mathcal{N}(i)$ stands for the neighborhood of $A_{i}$ in $G_{D}$. In what follows, we introduce an enhanced version of LLL, leveraging additional local structural information.
@@
(Quantitative Clique Lovász Local Lemma [21]) Let $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{n}\}$ be a set of events with dependency graph $G_{D}$ and let $\mathcal{K}=\{K_{1},K_{2},\ldots,K_{c}\}$ be a set of cliques in $G_{D}$ covering all the edges (not necessarily disjointly). If there exist a set of vectors $\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{c}\}$ from $(0,1)^{n}$ such that following conditions are satisfied:
@@
    $\forall i\in[n],\forall v$ such that $i\in K_{v}$:
@@
    In the variable framework, the running time of the Moser–Tardos algorithm is at most:
@@
III Upper Bound and Construction
@@ III-A Upper Bounds for Lifting Degree and Memory
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph $G_{base}=(V,E)$, with no parallel edges. The harmful weight factor $W_{\mathcal{H}}^{e}$ of an edge $e$ in the base graph with respect to the harmful structure set =$\mathcal{H}=\{H_{1},...,H_{k}\}$ is the number of harmful structures passing through the edge in the base graph that belong to the above set. Let $W_{\mathcal{H}}:=\text{Max}\{W_{\mathcal{H}}^{e}|e\in E(G_{base})\}$ be the maximum harmful weight factor.
@@
Consider a random cyclic $Z$-lift of a base bipartite graph $G$ with no parallel edges, and consider a harmful structure $H$ with $n_{b}$ fundamental cycles $\{c_{1},\dots,c_{n_{b}}\}$. Let $P_{H}^{l}$ denote the probability that $H$ is active after a random lifting. We then have
@@
The proof is obtained by combining [22, Theorem 2] with [23, Lemma 3]. ∎
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$. =$\mathcal{H}=\{H_{1},\ldots,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},\ldots,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading and lifting process, if
$\text{Max}\{P_{H_{1}},\ldots,P_{H_{k}}\}\leq\text{Max}\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{(|H|-1)^{|H|-1}}{(W_{\mathcal{H}}-1)|H|^{|H|}}}_{II}\}$
Here, $\Delta$ is the maximum degree of the dependency graph, and $|H|=\text{Max}\{|H_{1}|,\ldots,|H_{k}|\}$ is the maximum number of vertices in the harmful structures. Then
${P(\bigcap_{i\in[k]}\bar{H_{i}})}\geq\begin{cases}(1-\frac{2}{\Delta})^{||G||},&{\text{if}}\ I>II\\ {(1-\frac{W_{\mathcal{H}}}{(W_{\mathcal{H}}-1)|H|})^{\gamma\kappa},}&{\text{otherwise.}}\end{cases}$ (6)
@@
    The probability that avoidable harmful structures $H_{k}$ is active after the random edge-spreading and lifting process is $P_{H_{k}}=P_{H_{k}}^{l}\times P_{H_{k}}^{s}$.
@@
By utilizing the structural information of the dependency graph, we reconstruct the set of cliques $\{K_{1},K_{2},\ldots,K_{c}\}$ to obtain tighter bounds. Note that for each vertex in the dependency graph, many of its neighbors are interconnected, allowing us to group the vertex’s neighbors into fewer, larger cliques. For each edge in the $\gamma\times\kappa$ base graph, the set of harmful structures passing through it forms a clique in the dependency graph. The set of these cliques $\{K_{1},K_{2},...,K_{\gamma\kappa}\}$ clearly covers all edges in the dependency graph. In this case, the neighbors of each vertex in the dependency graph can be grouped into $|H|$ cliques. Assuming $\forall v\in[c],i\in[k]$, $x_{i,v}=x$, the conditions in Lemma 2 reduces to:
@@
Since $\Delta>|H|$ and $(|H|-1)|H|>\Delta$ always hold, we obtain:
@@
$P(H_{i})\leq\text{Max}\left\{\frac{(\Delta-1)^{\Delta-1}}{\Delta^{\Delta}},\frac{(|H|-1)^{|H|-1}}{(\Delta-1)|H|^{|H|}}\right\}$ (19)
@@
$\text{Max}\{P_{H_{1}},\ldots,P_{H_{k}}\}\leq\text{Max}\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{(|H|-1)^{|H|-1}}{(W_{\mathcal{H}}-1)|H|^{|H|}}}_{II}\}$
@@
${P(\bigcap_{i\in[k]}\bar{H_{i}})}\geq\begin{cases}(1-\frac{2}{\Delta})^{||G||},&{\text{if}}\ I>II\\ {(1-\frac{W_{\mathcal{H}}}{(W_{\mathcal{H}}-1)|H|})^{\gamma\kappa},}&{\text{otherwise.}}\end{cases}$ (20)
@@
$\frac{2m^{2}+4m+3}{3(m+1)^{3}Z}\leq\text{Max}\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{27}{256(\gamma\kappa-\gamma-\kappa)}}_{II}\}$ (22)
@@
When $\mathcal{H}=\mathcal{C}_{4}$, $|H|=4$, $|G|=\binom{\gamma}{2}\binom{\kappa}{2}$, $||G||\leq\frac{|G|\Delta}{2}$, $W_{\mathcal{C}_{4}}=\binom{\gamma-1}{1}\binom{\kappa-1}{1}$. ∎

### v2

@@ I Introduction
Spatially-coupled low-density parity-check (SC-LDPC) codes have attracted much attention due to their threshold saturation property [1, 2, 3] and the capacity-achieving performance over general binary memoryless channels [4]. In addition, the highly regular structure makes it possible to decode using low-latency windowed decoding (WD) [5], [6]. Research on SC-LDPC codes can be traced back to the convolutional LDPC codes introduced by Felström and Zigangirov [7].
In practical data storage systems, the frame error rate (FER) requirement is below $10^{-12}$ [8, 9, 10]. Therefore, it is crucial to construct SC-LDPC codes with good bit-error rate (BER) performance in both waterfall and error floor regions under belief-propagation (BP) decoders. In the low-SNR region, the performance of SC-LDPC codes is mainly determined by their asymptotic properties (such as decoding threshold) [4], and density evolution (DE) techniques can be used to analyze the decoding threshold of the code. In the high-SNR region, the performance of SC-LDPC codes is determined by finite-length characteristics associated with harmful topological structures (e.g., cycles and trapping sets) [9], [10].
@@
In this paper, we consider the non-uniform protograph-based SC-LDPC codes in the same probability metric as [14], and provide mathematical support for the construction of SC-LDPC codes through probabilistic combinatorial methods. In the extended version [15], we provide complete proofs of this work, along with further discussion of the new bounds and constructions.
@@ II-A QC-SC-LDPC Codes
In this section, we introduce the construction process of QC-SC-LDPC codes using QC-LDPC codes as underlying block codes. The QC structure not only enables efficient encoding and decoding implementation to SC-LDPC codes [16], [17], but also makes the constructed codes exhibit anytime-reliable properties [18], [19].
We consider the QC-SC-LDPC code of Type-I. First, we start with an all-one $(\gamma,\kappa)$ base matrix $\mathbf{H}$, which corresponds to a fully connected bipartite base graph. As the name suggests, SC-LDPC codes are generally constructed by coupling a series of identical block codes into a chain. This operation has different names in different perspectives, but the most widely intuitive is the matrix perspective (edge-spreading), which is seen as dividing the edges of the base matrix $\mathbf{H}$ into $m+1$ component matrices $\{\mathbf{H}_{0},...,\mathbf{H}_{m}\}$ of the same size, such that $\mathbf{H}=\sum_{l=0}^{m}\mathbf{H}_{l}$, where $m$ is memory. We record the exact way in which the edge spreads in the $(\gamma,\kappa)$ partition matrix $\mathbf{P}$, which means that $\mathbf{P}(i,j)=k,~0\leq k\leq m,~i\in[\gamma],~j\in[\kappa]$, if and only if $\mathbf{H}_{k}(i,j)=1,~\mathbf{H}_{l}(i,j)=0,~\forall l\neq k,~0\leq l\leq m$. Generally, the base matrices and edge-spreading patterns may differ among uncoupled codes, like those constructed via the random spreading method for the proof of universality. However, considering the implementation complexity in practical systems, the base matrices and the edge-spreading patterns of all uncoupled codes are assumed to be identical in this paper. By concatenating these component matrices vertically from $\mathbf{H}_{0}$ to $\mathbf{H}_{m}$, a replica is obtained. By concatenating $L$ of these identical replicas horizontally in space, we get the protograph matrix $\mathbf{H}_{SC}^{P}$ of the SC-LDPC code, where $L$ is called the coupling length.
To construct QC-SC-LDPC codes, after the edge-spreading process, we perform the lifting operation on the protograph. The lifting for every uncoupled code is identical, which facilitates the practical implementation of the windowed decoder. Let $\mathbf{L}$ be a $\gamma\times\kappa$ matrix called the lifting matrix. $\sigma$ denotes the $Z\times Z$ circulant matrix obtained by cyclically shifting the columns of an identity matrix one unit to the left. For all $i\in[\gamma],~j\in[\kappa]$, if $\mathbf{L}(i,j)=x$ where $x\in\{0,\dots,Z-1\}$, then replace the ones in $\mathbf{H}_{SC}^{P}$ corresponding to $\mathbf{H}(i,j)$ with the $\sigma^{x}$ matrix, and replace the remaining zeros with $Z\times Z$ zero matrices. Under this notation, the parameters of a QC-SC-LDPC code $\mathbf{H}_{SC}$ are described by $(\gamma,\kappa,m,Z,L)$.
@@ II-B Major Harmful Structures
Despite the fact that QC-SC-LDPC codes offer improved performance compared to their underlying block codes, they still experience performance degradation at high signal-to-noise ratios (SNRs) over the additive white Gaussian noise (AWGN) channel under iterative decoding, resulting in an error floor phenomenon. This error floor is primarily caused by small structures in the code’s Tanner graph, known as trapping sets (TSs), which, in turn, consist of small cycles. We review the key definitions:
@@
Let $\mathcal{C}_{2g}$ be the set of cycle-2g candidates in the base matrix, where $g\in\mathbb{N},~g\geq 2$. Denote $c_{2g}\in\mathcal{C}_{2g}$ by $(j_{1},i_{1},j_{2},i_{2},\dots,j_{g},i_{g})$, where $(i_{k},j_{k})$ and $(i_{k},j_{k+1})$ for $1\leq k\leq g$, with $j_{g+1}=j_{1}$, are nodes of $c_{2g}$ in $\mathbf{H}$, $\mathbf{P}$, and $\mathbf{L}$. Then $c_{2g}$ becomes a cycle candidate in the protograph after the edge-spreading operation if and only if the following condition holds [20]:
@@
To reduce the complexity of optimizing QC-SC-LDPC codes, the process is typically divided into two stages. First, matrix $\mathbf{P}$ is optimized to minimize cycle candidates in the protograph. This sets the stage for optimizing matrix $\mathbf{L}$, aiming to further reduce cycles in the Tanner graph. This two-step method efficiently streamlines optimization and reduces cycles, as shown in [12], [14].
In the remainder of this paper, we focus on QC-SC-LDPC codes for the AWGN channel, where the dominant harmful structures are the low-weight absorbing sets (ASs) [12]. The ASs are defined as follows:
@@ II-C Lovász Local Lemma
The Lovász Local Lemma (LLL), introduced by Erdős and Lovász in 1975, has become one of the most important probabilistic methods. LLL provides sufficient conditions under which a set of undesirable events $\mathcal{A}$ in a probability space $\Omega$ can be simultaneously avoided, meaning the conditions for $P(\cap_{A\in\mathcal{A}}\bar{A})>0$ to hold. The relationships between these bad events are depicted by an undirected graph called the dependency graph $G_{D}=([n],E)$ with $|G|:=n$ and $\|G\|:=|E|$, where each bad event corresponds to a vertex in the dependency graph. An event $A_{i}$ is independent of $\{A_{j}:j\neq i,j\notin\mathcal{N}(i)\}$, where $\mathcal{N}(i)$ stands for the neighborhood of $A_{i}$ in $G_{D}$. The scope of an event $A$, denoted $sc(A)$, is the minimal set of random variables on which $A$ depends. In what follows, we introduce an enhanced version of LLL, leveraging additional local structural information.
@@
(Quantitative Clique Lovász Local Lemma [22]) Let $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{n}\}$ be a set of events with dependency graph $G_{D}$ and let $\mathcal{K}=\{K_{1},K_{2},\ldots,K_{c}\}$ be a set of cliques in $G_{D}$ covering all the edges (not necessarily disjointly). If there exists a set of vectors $\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{c}\}$ from $(0,1)^{n}$ such that following conditions are satisfied:
@@
    $\forall i\in[n],~\forall v$ such that $i\in K_{v}$:
@@
    In the variable framework, the running time of the Moser–Tardos algorithm ( Algorithm 1 ) is at most:
@@
III Upper Bounds and Constructions
@@ III-A Upper Bounds for Lifting Degree and Memory
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph $G_{B}=(V,E)$, with no parallel edges. The harmful weight factor $W_{\mathcal{H}}^{e}$ of an edge $e$ in the base graph with respect to the harmful structure set $\mathcal{H}=\{H_{1},...,H_{k}\}$ is the number of harmful structures passing through the edge in the base graph that belong to the above set. Let $W_{\mathcal{H}}:=\max\{W_{\mathcal{H}}^{e}|e\in E(G_{B})\}$ be the maximum harmful weight factor.
@@
Consider a random cyclic $Z$-lifting of a base bipartite graph $G$ with no parallel edges, and consider a harmful structure $H$ with $n_{b}$ fundamental cycles $\{c_{1},\dots,c_{n_{b}}\}$. Let $P_{H}^{l}$ denote the probability that $H$ is active after a random lifting. We then have
@@
The proof is obtained by combining [23, Theorem 2] with [24, Lemma 3]. ∎
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$. $\mathcal{H}=\{H_{1},\ldots,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},\ldots,P_{H_{k}}\}=\{P(H_{1}),\ldots,P(H_{k})\}$ are the probabilities that they are active after the random edge-spreading and lifting process, if
$\max\{P_{H_{1}},\ldots,P_{H_{k}}\}\leq\max\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{(|H|-1)^{|H|-1}}{(W_{\mathcal{H}}-1)|H|^{|H|}}}_{II}\}$
Here, $\Delta$ is the maximum degree of the dependency graph, and $|H|=\max\{|H_{1}|,\ldots,|H_{k}|\}$ is the maximum number of vertices in the harmful structures. Then
${P(\bigcap_{i\in[k]}\bar{H_{i}})}\geq\begin{cases}(1-\frac{2}{\Delta})^{\|G\|},&{\text{if}}\ I>II\\ {(1-\frac{W_{\mathcal{H}}}{(W_{\mathcal{H}}-1)|H|})^{\gamma\kappa},}&{\text{otherwise.}}\end{cases}$ (6)
@@
    The probability that avoidable harmful structures $H_{i}$ is active after the random edge-spreading and lifting process is $P(H_{i})=P_{H_{i}}^{l}\times P_{H_{i}}^{s}$.
@@
By utilizing the structural information of the dependency graph, we reconstruct the set of cliques $\{K_{1},K_{2},\ldots,K_{c}\}$ to obtain tighter bounds. Note that for each vertex in the dependency graph, many of its neighbors are interconnected, allowing us to group the vertex’s neighbors into fewer, larger cliques. For each edge in the $\gamma\times\kappa$ base graph, the set of harmful structures passing through it forms a clique in the dependency graph. The set of these cliques $\{K_{1},K_{2},...,K_{\gamma\kappa}\}$ clearly covers all edges in the dependency graph. In this case, the neighbors of each vertex in the dependency graph can be grouped into $|H|$ cliques. Assuming $\forall v\in[c],~i\in[k]$, $x_{i,v}=x$, the conditions in Lemma 2 reduces to:
@@
Since $\Delta>|H|$ and $(W_{\mathcal{H}}-1)|H|>\Delta$ always hold, we obtain:
@@
$P(H_{i})\leq\max\left\{\frac{(\Delta-1)^{\Delta-1}}{\Delta^{\Delta}},\frac{(|H|-1)^{|H|-1}}{(W_{\mathcal{H}}-1)|H|^{|H|}}\right\}$ (19)
@@
$\max\{P_{H_{1}},\ldots,P_{H_{k}}\}\leq\max\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{(|H|-1)^{|H|-1}}{(W_{\mathcal{H}}-1)|H|^{|H|}}}_{II}\}$
@@
[…]
@@
$\frac{2m^{2}+4m+3}{3(m+1)^{3}Z}\leq\max\{\underbrace{\frac{(\Delta-1)^{\Delta-1}}{(\Delta)^{\Delta}}}_{I},\underbrace{\frac{27}{256(\gamma\kappa-\gamma-\kappa)}}_{II}\}$ (22)
@@
When $\mathcal{H}=\mathcal{C}_{4}$, $|H|=4$, $|G|=\binom{\gamma}{2}\binom{\kappa}{2}$, $G||\leq\frac{|G|\Delta}{2}$, $W_{\mathcal{C}_{4}}=\binom{\gamma-1}{1}\binom{\kappa-1}{1}$. ∎